# Understanding Addition in Transformers

### v1

@@ Abstract
Understanding the inner workings of machine learning models like Transformers is vital for their safe and ethical use. This paper presents an in-depth analysis of a one-layer Transformer model trained for integer addition. We reveal that the model divides the task into parallel, digit-specific streams and employs distinct algorithms for different digit positions. Our study also finds that the model starts calculations late but executes them rapidly. A rare use case with high loss is identified and explained. Overall the model’s algorithm is explained in detail. These findings are validated through rigorous testing and mathematical modeling, contributing to the broader works in Mechanistic Interpretability, AI safety, and alignment. Our approach opens the door for analyzing more complex tasks and multi-layer Transformer models.
@@ 1 Introduction
In this work, we offer an in-depth analysis of a one-layer Transformer model performing integer addition. We show that the model separates the addition task into independent digit-specific streams of work, which are computed in parallel. Different algorithms are employed for predicting the first, middle, and last digits of the answer. The model’s behavior is influenced by the compact nature of the task and the specific format in which the question is presented. Despite having the opportunity to begin calculations early, the model actually starts later. The calculations are performed in a time-dense manner, enabling the model to add two 5-digit numbers to produce a 6-digit answer in just 6 layers (See Fig. 1). A rare use case with high loss was predicted by analysis and proved to exist via experimentation. Our findings shed light on understanding and interpreting transformers. These insights may also have implications for AI safety and alignment.
As evidenced in appendices 11 and 12, our results demonstrate the model’s unique approach to integer addition across various digit lengths. The theoretical framework we develop provides a mathematical justification for the model’s behavior, substantiating our empirical observations and offering a foundation for future work.
@@
Figure 1: Illustration of the transformer model’s attention pattern when adding two 5-digit integers. The model attends to digit pairs sequentially from left to right, resulting in a "double staircase" pattern across layers. A: The 5 digit question is revealed token by token. The “10s of thousands" digit is revealed first. B: From the “=" token, the model attention heads focus on successive pairs of digits, giving a “double staircase" attention pattern. C: The 3 heads are time-offset from each other by 1 token such that, in each layer, data from 3 tokens is available. D: To calculate A3, the 3 heads do independent simple calculations on D3, D2 and D1. The results are combined by the MLP layer using trigrams. A3 is calculated one token before it is needed. This approach applies to all answer digits, with the first and last digits using slight variations of the approach.
@@ 2 Background
We focus on a single-layer transformer model with a vocabulary of size $V$ containing the symbols “0” to “9", “+" and “=". The model converts the human readable input (e.g. “12345+67890=") into an input sequence $(\texttt{x}_{1},\ldots,\texttt{x}_{p})$ where each $\texttt{x}_{i}\in\{1,\ldots,V\}$. Tokens are mapped to $d\_e$-dimensional embeddings by selecting the $\texttt{x}_{i}$-th column of $E\in\mathbb{R}^{d_{e}\times V}$. The model processes the input tokens, using a mechanism called “self-attention". Each input token is passed through a self-attention mechanism that calculates weighted relationships between all input tokens - capturing the importance of each token relative to others. The model then aggregates these weighted representations to produce contextually enriched representations for each token. These enriched representations are subsequently fed through feedforward neural networks (i.e. an MLP) to refine their information. Finally, the output tokens are generated based on the refined representations, and converted back to human readable format using the vocabulary (e.g. “12345+67890=80235").
@@ 3 Related Work
Interpreting and reverse engineering neural networks and transformers to find meaningful circuits has been an area of active research. Olah et al. [2020a] argued that by studying the connections between neurons and their weights, we can find meaningful algorithms (aka Circuits) in a “vision” neural network.
Elhage et al. [2021] extended this approach to transformers, conceptualizing their operation in a mathematical framework that allows significant understanding of how transformer operate internally. Tools (such as explained in Foote et al. [2023], Conmy et al. [2023b]) use this framework to semi-automate some aspects of reverse engineering. Nanda et al. [2023] reverse-engineered modular addition (e.g. 5 + 7 mod 10 = 2) showing the model used discrete Fourier transforms and trigonometric identities to convert addition to rotation about a circle.
Further, Nanda and Lieberum [2022] have argued models comprise multiple circuits, evidenced by distinct training loss curves per subtask. However, the underlying circuits remained uninvestigated. This work extends such efforts by detailing identified circuits.
From a circuit analysis lens, studies like extract graphical circuit representations and analyze component interactions. To enable analysis and interpretability, techniques in works like symbolically reverse-engineer networks by recovering computational expressions.
Research including Seth [2005] advocate analyzing networks causally, introducing structural causal models to infer mechanisms. Examinations of sequence models like have analyzed the emergence and interaction of modular components during training.
Evolutionary perspectives such as Miikkulainen [2021] elucidate how selection pressures shape hierarchical representations.
Information bottleneck analyses including relate bottlenecks to abstractions and modularization arising from forced compression. Surveys like overview techniques to disentangle explanatory factors into separate latent dimensions.
Novel objectives proposed in works like Conmy et al. [2023a] improve interpretability by encouraging modularity and disentanglement.
@@ 4 Methodology
The traditional human addition process first sums the units before moving on to higher value digits. This is the simplest process but relies on being able to choose the order to process the digits in. Transformer models always process text from left to right. So the model sees the higher value digits (e.g. thousands) of the question before the lower value digits (e.g. units). It doesn’t use the traditional process.
A key component of addition is the need to sum each digit in the first number with the corresponding digit in the second number. Transformer models contain “attention heads" and they are the only computational sub-component of a model that can move information between positions (aka digits or tokens). Visualising which token(s) each attention head focussed on in each layer of the calculation provided insights. While our model works with 2, 3 or 4 attention heads, 3 attention heads gives the most easily interpreted attention patterns. Fig. 3 shows the attention pattern for a single 5 digit addition calculation using 3 attention heads. Appendix 12 shows the same pattern for 10 and 15 digit addition. Appendix 12 shows the pattern with 2 or 4 attention heads.
Figure 3: The attention pattern, for a model with 3 attention heads, performing a single 5 digit addition. The pattern is 18 by 18 squares (as 54321+77779=132100 is 18 tokens). Time proceeds vertically downwards, with one additional token being revealed horizontally at each layer, giving the overall triangle shape. After the question is fully revealed (at layer 11), each head starts attending to pairs of question digits from left to right (i.e. high-value digits before lower-value digits) giving the “double staircase" shape. The three heads attend to a given digit pair in three different layers, giving a time ordering of heads.
@@ 5 Mathematical Framework
Figure 4: The mathematical framework (our method) predicts that during training, tasks are learnt for each digit independently, progressively increasing per digit accuracy (i.e. decreasing loss) shown as percentages). Mathematical rules cause dependencies between digits, giving an predicted ordering for perfect (i.e. zero loss) addition. The chain of blue squares relate to questions like 99999 + 00001 = 100000 where the MC1 in digit 0 causes US9 cascades through multiple other digits.
@@ 6 Training Analysis
The Digit 0 calculation is the least interesting as it only uses BA (not UC1 or US9 ). Once discovered, Digit 0 always quickly refines to have the lowest loss and least noise (as expected). Refer Appendix 11 for graphs.
@@
Figure 6: Training loss for digit 3 showing that, in Phase 2, the refining of Use Carry 1 lags behind Base Add . Base Add and Use Carry 1 are refined separately and have separate calculation algorithms.
@@ 7 Prediction Analysis
During model prediction, we overrode (mean ablated) the model memory (residual stream) at each layer, and confirmed that the addition algorithm does not use any data generated in layers 0 to 10 inclusive. In these layers the model has not yet seen the full question and every digit in the question is independent of every other digit, making accurate answer prediction infeasible. The model also does not use the last (17th) layer. Therefore, the addition is started and completed in 6 layers (11 to 16). Further (ablation) experiments confirmed that the A0 to A4 answers are calculated one layer before being revealed.
@@
Figure 8: A: For A3, the addition algorithm must combine information from digits 3, 2 and 1. B: The 1st head calculates MC1 on digit 1. C: The 2nd head calculates MC1 and MS9 (at most one of which can be true) on digit 2. D: The 3rd head calculates Base Add on digit 3. E: The MLP layer uses trigrams to combine the information from the 3 heads to give the final answer A3, one layer (token) before it is output.
@@
The A3 algorithm could also successfully be applied to A4. But the Digit 4 training curve is better (faster) than the middle digits. The attention patterns shows that for A4, the model is using all the heads in layer 11 (the “=" token) when the A3 algorithm doesn’t require this. Uniquely, A4 utilises more “compute" than is available to A3, A2, A1 or A0. We assume the model uses this advantage to implement a faster-training and lower-loss algorithm for A5 and A4. We haven’t worked out the details of this.
Mean ablating the 1st or 2nd head slightly increased the average loss for BA questions from 0.05 to 0.08, whereas ablating the 3rd head substantially increased the loss to 3.7, confirming that the 3rd head is doing the BA task.
@@
Despite being feasible, the model does not calculate the task MC1 in layers 7 to 11. Instead it completes each digit calculation in 1 layer, possibly because there are training optimisation benefits in generating a “compact" algorithm.

### v5

@@ Abstract
Understanding the inner workings of machine learning models like Transformers is vital for their safe and ethical use. This paper presents an in-depth analysis of a one-layer Transformer model trained for $n$-digit integer addition. We reveal that the model divides the task into parallel, digit-specific streams and employs distinct algorithms for different digit positions. Our study also finds that the model starts calculations late but executes them rapidly. A rare use case with high loss is identified and explained. Overall the model’s algorithm is explained in detail. These findings are validated through rigorous testing and mathematical modeling, contributing to the broader works in Mechanistic Interpretability, AI safety, and alignment. Our approach opens the door for analyzing more complex tasks and multi-layer Transformer models. Code for all experiments is available at https://github.com/apartresearch/conceptual-interp
@@ 1 Introduction
In this work, we offer an in-depth analysis of a one-layer Transformer model performing $n$-digit integer addition. We show that the model separates the addition task into independent digit-specific streams of work, which are computed in parallel. Different algorithms are employed for predicting the first, middle, and last digits of the answer. The model’s behavior is influenced by the compact nature of the task and the specific format in which the question is presented. Despite having the opportunity to begin calculations early, the model actually starts later. The calculations are performed in a time-dense manner, enabling the model to add two 5-digit numbers to produce a 6-digit answer in just 6 steps (See Fig. 1). A rare use case with high loss was predicted by analysis and proved to exist via experimentation. Our findings shed light on understanding and interpreting transformers. These insights may also have implications for AI safety and alignment.
Our results demonstrate the model’s unique approach applies to integer addition across various digit lengths (Refer Appendixes B and C). Our theoretical framework provides a mathematical justification for the model’s behavior, substantiating our empirical observations and offering a foundation for future work.
@@
Figure 1: Illustration of the transformer model’s attention pattern when adding two 5-digit integers. The model attends to digit pairs sequentially from left to right, resulting in a "double staircase" pattern across rows. A: The 5 digit question is revealed token by token. The “10s of thousands" digit is revealed first. B: From the “=" token, the model attention heads focus on successive pairs of digits, giving a “double staircase" attention pattern. C: The 3 heads are time-offset from each other by 1 token such that, in each row, data from 3 tokens is available. D: To calculate A3, the 3 heads do independent simple calculations on D3, D2 and D1. The results are combined by the MLP layer using trigrams. A3 is calculated one token before it is needed. This approach applies to all answer digits, with the first and last digits using slight variations of the approach.
@@ 2 Background
We focus on a single-layer transformer model with a vocabulary of size $V$ containing the symbols “0” to “9", “+" and “=". The model converts the human readable input (e.g. “12345+67890=") into an input sequence $(\texttt{x}_{1},\ldots,\texttt{x}_{p})$ where each $\texttt{x}_{i}\in\{1,\ldots,V\}$. Tokens are mapped to de dimensional embeddings by selecting the $\texttt{x}_{i}$-th column of $E\in\mathbb{R}^{d_{e}\times V}$. The model processes the input tokens, using a mechanism called “self-attention". Each input token is passed through a self-attention mechanism that calculates weighted relationships between all input tokens - capturing the importance of each token relative to others. The model then aggregates these weighted representations to produce contextually enriched representations for each token. These enriched representations are subsequently fed through feedforward neural networks (i.e. an MLP) to refine their information. Finally, the output tokens are generated based on the refined representations, and converted back to human readable format using the vocabulary (e.g. “12345+67890=80235").
@@ 3 Related Work
Interpreting and reverse engineering neural networks and transformers to find meaningful circuits has been an area of active research. Olah et al. [2020a] argued that by studying the connections between neurons and their weights, we can find meaningful algorithms (aka Circuits) in a “vision” neural network. Elhage et al. [2021] extended this approach to transformers, conceptualizing their operation in a mathematical framework that allows significant understanding of how transformer operate internally. Various tools [Foote et al., 2023, Conmy et al., 2023b] use this framework to semi-automate some aspects of reverse engineering. Nanda et al. [2023] reverse-engineered modular addition (e.g. 5 + 7 mod 10 = 2) showing the model used discrete Fourier transforms and trigonometric identities to convert modular addition to rotation about a circle.
Nanda and Lieberum [2022] have argued models comprise multiple circuits. They gave examples, including the distinct training loss curve per answer digit in 5-digit integer addition, but did not identify the underlying circuits. This work investigates and explains the circuits in n-digit integer addition.
@@
Circuit analysis can extract graphical circuit representations and analyze component interactions [Bau et al., 2017]. To enable analysis and interpretability, techniques in works like Petersen et al. [2021] symbolically reverse-engineer networks by recovering computational expressions. Research including Seth [2005] advocate analyzing networks causally, introducing structural causal models to infer mechanisms. Examinations of sequence models like Petersen et al. [2021] have analyzed the emergence and interaction of modular components during training. Evolutionary perspectives such as Miikkulainen [2021] elucidate how selection pressures shape hierarchical representations.
Information bottleneck analyses including Kawaguchi et al. [2023] relate bottlenecks to abstractions and modularization arising from forced compression. Surveys like Carbonneau et al. [2022] overview techniques to disentangle explanatory factors into separate latent dimensions. Novel objectives proposed in works like Conmy et al. [2023a] improve interpretability by encouraging modularity and disentanglement.
@@ 4 Methodology
The traditional human addition process first sums the units before moving on to higher value digits. This is the simplest process but relies on being able to choose the order to process the digits in. This autoregressive transformer model processes text from left to right. So the model sees the higher value digits (e.g. thousands) of the question before the lower value digits (e.g. units). It doesn’t use the traditional process.
A key component of addition is the need to sum each digit in the first number with the corresponding digit in the second number. Transformer models contain “attention heads" and they are the only computational sub-component of a model that can move information between positions (aka digits or tokens). Visualising which token(s) each attention head focussed on in each row of the calculation provided insights. While our model works with 2, 3 or 4 attention heads, 3 attention heads gives the most easily interpreted attention patterns. Fig. 3 shows the attention pattern for a single 5 digit addition calculation using 3 attention heads. Appendix C shows the same pattern for 10 and 15 digit addition. Appendix C shows the pattern with 2 or 4 attention heads.
Figure 3: The attention pattern, for a model with 3 attention heads, performing a single 5 digit addition. The pattern is 18 by 18 squares (as 54321+77779=132100 is 18 tokens). Time proceeds vertically downwards, with one additional token being revealed horizontally at each row, giving the overall triangle shape. After the question is fully revealed (at row 11), each head starts attending to pairs of question digits from left to right (i.e. high-value digits before lower-value digits) giving the “double staircase" shape. The three heads attend to a given digit pair in three different rows, giving a time ordering of heads.
@@ 5 Mathematical Framework
Figure 4: The mathematical framework (our method) predicts that during training, tasks are learnt for each digit independently, progressively increasing per digit accuracy (i.e. decreasing loss) shown as percentages. Mathematical rules cause dependencies between digits, giving an predicted ordering for perfect (i.e. zero loss) addition. The chain of blue squares relate to questions like 99999 + 00001 = 100000 where the MC1 in digit 0 causes US9 cascades through multiple other digits.
We use this mathematical framework solely for analysis to gain insights. The model training and all loss calculations are completely independent of this mathematical framework.
@@ 6 Training Analysis
The Digit 0 calculation is the least interesting as it only uses BA (not UC1 or US9 ). Once discovered, Digit 0 always quickly refines to have the lowest loss and least noise (as expected). (Graphs in Appendix B.)
@@
Figure 6: Training loss for digit 3 showing that, in Phase 2, the refining of Use Carry 1 lags behind Base Add . Base Add and Use Carry 1 are refined separately and have separate calculation algorithms. The 3 phases seem to correspond to “memorisation”, “algorithm discovery” and “clean-up”.
@@ 7 Prediction Analysis
During model prediction, we overrode (mean ablated) the model memory (residual stream) at each row, and confirmed that the addition algorithm does not use any data generated in rows 0 to 10 inclusive. In these rows the model has not yet seen the full question and every digit in the question is independent of every other digit, making accurate answer prediction infeasible. The model also does not use the last (17th) row. Therefore, the addition is started and completed in 6 rows (11 to 16). Further (ablation) experiments confirmed that the A0 to A4 answers are calculated one row before being revealed. (Details in Appendix H.)
@@
Figure 8: A: For A3, the addition algorithm must combine information from digits 3, 2 and 1. B: The 1st head calculates MC1 on digit 1. C: The 2nd head calculates MC1 and MS9 (at most one of which can be true) on digit 2. D: The 3rd head calculates Base Add on digit 3. E: The MLP layer uses trigrams to combine the information from the 3 heads to give the final answer A3, one row before it is output. Appendix G shows this algorithm as pseudocode.
@@
The A3 algorithm could also successfully be applied to A4. But the Digit 4 training curve is better (faster) than the middle digits. The attention patterns shows that for A4, the model is using all the heads in row 11 (the “=" token) when the A3 algorithm doesn’t require this. Uniquely, A4 utilises more “compute" than is available to A3, A2, A1 or A0. We assume the model uses this advantage to implement a faster-training and lower-loss algorithm for A5 and A4. We haven’t worked out the details of this.
Mean ablating the 1st or 2nd head slightly increased the average loss for BA questions from 0.05 to 0.08, whereas ablating the 3rd head substantially increased the loss to 3.7, confirming that the 3rd head is doing the BA task. (Details in Appendix H.)
@@
Despite being feasible, the model does not calculate the task MC1 in rows 7 to 11. Instead it completes each digit calculation in 1 row, possibly because there are training optimisation benefits in generating a “compact" algorithm.
@@ 9 Reproducibility Statement
Appendix A Appendix - Number of Attention Heads
The model can be successfully trained with 2, 3 or 4 attention heads. However as described in (Fig. 10), 3 attention heads is more useful for human understanding than 2 or 4 heads.
Figure 9: For 5 digit addition, using 4 attention heads gives an attention pattern where the H1 and H2 staircases overlap perfectly. Ablating one of H1 or H2 increases loss. The similarity in H1 and H2’s attention patterns suggests they are “splitting" a single logical task. Spliting is feasible as Elhage et al. [2021] says attention heads are independent and additive.
Figure 10: For 5 digit addition, using 3 attention heads gives the best separation with the heads having distinct, non-overlapping attention pattern. Ablating any head increases the model loss, showing all 3 heads are useful.
Figure 11: For 5 digit addition, using 2 attention heads works, but the model attends to multiple token pairs in a single head, suggesting that multiple tasks are being packed into a single head, which makes it harder to understand.
Appendix B Appendix - Training Loss by Number of Digits
The model can be successfully trained with 2, 5, 10, 15, etc digits. As expected, for a given loss threshhold, 15 digit addition takes longer to train than 10 digit addition, which takes longer to train than 5 digit addition.
Figure 12: The per-digit training loss curves for 5 digit addition
Figure 13: The per-digit training loss curves for 10 digit addition.
Figure 14: The per-digit training loss curves for 15 digit addition.
Appendix C Appendix - Attention Patterns by Number of Digits
The model can be successfully trained with 2, 5, 10, 15, etc digits. The attention patterns for theses cases show similarities that aid in human understanding.
Figure 15: For 5 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the heads are attending to different tokens in each row.
Figure 16: For 10 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the heads are attending to different tokens in each row.
Figure 17: For 15 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the heads are attending to different tokens in each row.
Appendix D Appendix - Model Configuration
Experimentation was done in a CoLab notebook:
  • It will be loaded to github post publication.
  • It runs on a T4 GPU with each experiment taking a few minutes to run.
  • The key parameters (which can all be experimented with) are:
    1. 1.
      n_layers = 1; This is a one layer Transformer
    2. 2.
      n_heads = 3; There are 3 attention heads
    3. 3.
      n_digits = 5; Number of digits in the addition question
  • It uses a new batch of data each training step (aka Infinite Training Data) to minimise memorisation.
  • During a training run the model processes about 1.5 million training datums. For the 5 digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
  • Because US9 cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare, the data generator was enhanced to increase the likelihood of these cases turning up in the training data.
The CoLab notebook can be downloaded from https://github.com/apartresearch/conceptual-interp
Appendix E Appendix - Solving a jigsaw using many people
Here is useful analogy for changing how we approach a problem to more closely resemble how a Transformer would; a jigsaw puzzle. A single person could solve it using a combination of meta knowledge of the problem (placing edge pieces first), categorisation of resources (putting like-coloured pieces into piles), and an understanding of the expected outcome (looking at the picture on the box).
But if instead we had one person for each piece in the puzzle, who only knew their piece, and could only place it once another piece that it fit had already been placed, but couldn’t talk to the other people, and did not know the expected overall picture, the strategy for solving the jigsaw changes dramatically.
When they start solving the jigsaw, the 4 people holding corner pieces place them. Then 8 people holding corner-adjacent edge pieces can place them. The process continues, until the last piece is placed near the middle of the jigsaw.
We posit that this approach parallels how transformer models work. There is no pre-agreed overall strategy or communication or co-ordination between people (circuits) - just some “rules of the game" to obey. The people think independently and take actions in parallel. The tasks are implicitly time ordered by the game rules.
Appendix F Appendix - Use Sum 9 Training Loss Graph
Fig. 18 shows a training loss curve for just BA and UC1 . The model has high loss on Use Sum 9 cascades shown as high variability in the loss graph (See Fig. 19).
Figure 18: Training loss for digit 3 showing that, in Phase 2, the refining of Use Carry 1 lags behind Base Add . This supports the claim that Base Add and Use Carry 1 are refined separately and have separate calculation algorithms.
Figure 19: Training loss for digit 3 for the Base Add , Use Carry 1 and Use Sum 9 tasks, showing the model struggles to reduce loss on Use Sum 9 compared to Base Add and Use Carry 1 .
Appendix G Appendix - Model Algorithm as Pseudocode
This section reproduces the algorithm summarised in Fig. 8 as pseudocode. This code calculates the largest digit first, handling simple (but not cascading) UseSum9 cases, and returning the answer.
Note that the pseudocode does not retain any information between passes through the for loop - this corresponds to each digit being calculated independently of the other digits.
Algorithm 1 n-digit integer addition algorithm
1:function CalculateAnswerDigits($n\ digits,q1,q2$)
2:   $\text{answer}\leftarrow 0$triangleright Initialize the answer to zero
3:   for all $i=0,\dots,n\ digits-1$ do $\triangleright$ Loop over each digit
4:      $\text{pos}\leftarrow n\ digits-i-1$triangleright Current position from the right
[…]
9:         if $q1_{\text{prev prev pos}}+q2_{\text{prev prev pos}}\geq 10$ then
[…]
11:         end if
12:      end if
[…]
15:         if $q1_{\text{prev pos}}+q2_{\text{prev pos}}\geq 10$ then
[…]
17:         end if
18:      end if
[…]
21:         if $q1_{\text{prev pos}}+q2_{\text{prev pos}}==9$ then
[…]
23:         end if
24:      end if
[…]
30:            end if
31:         end if
32:      end if
[…]
34:      $\text{digitanswer}\leftarrow\textsc{MODULUS}(\text{digitanswer},10)$triangleright Correct the current digit if it’s $\geq 10$
35:      $\text{answer}\leftarrow\text{digitanswer}+\text{answer}\times 10$triangleright Concatenate the current digit to the answer
36:   end forreturn answer $\triangleright$ Return the final calculated answer
37:end function
Appendix H Appendix - The Loss function and loss measures
The loss function is simple:
  • Per Digit Loss: For “per digit” graphs and analysis, for a given answer digit, the loss used is negative log likelihood.
  • All Digits Loss: For “all answer digits” graphs and analysis, the loss used is the mean of the “per digit” loss across all the answer digits.
The final training loss varies with the number of digits in the question as shown in Tab. 1.
 Size of question Final training loss Example question
 5 digit addition 0.009
11111 + 22222 = 033333
10 digit addition 0.011
1111111111 + 2222222222 = 03333333333
15 digit addition 0.031
111111111111111 + 222222222222222 = 0333333333333333
 Overall Average 0.017 General Performance
Table 1: The final training loss after 5000 training epochs, each containing 64 questions, using All Digits Loss, for different size addition models.
The final training loss for each digit in the question varies as shown in Tab. 2.
 Digit index Training loss AKA Example of digit
 5 <0.001 A5 11111+22222=033333
4 <0.001 A4 11111+22222=033333
3 0.003 A3 11111+22222=033333
2 0.008 A2 11111+22222=033333
1 0.046 A1 11111+22222=033333
0 0.001 A0 11111+22222=033333
 Overall Average 0.010
Table 2: For 5-digit addition, final training loss per digit.
Using our mathematics framework, we categorize each training question as one of following:
  • BA : BaseAdd questions only require “base add” calculations to get the correct answer.
  • MC1 : MakeCarry1 questions require “use carry 1” and “base add” calculations to get the correct answer.
  • US9 : UseSum9 questions require “use sum 9”, “use carry 1” and “base add” calculations to get the correct answer.
This allows us to measure the loss per category as shown in Tab. 3. The frequency of these question types differ significantly ( BA = 61%, MC1 =33%, US9 =6% ) so the final training loss values, at this level, are not very informative.
 Question type Training loss Aka Example of digit
 BA 0.021 Base Add 11111+22222=033333
MC1 0.001 Make Carry 1 11811+22222=034033
US9 <0.001 Use Sum 9 17811+22222=040033
 Overall Average 0.011 General Performance
Table 3: For 5 digit addition, the final training loss per question category.
After training, we used the model to give answers to questions. To understand whether the calculations at token n are important, we look at the impact on loss of ablating all attention heads at that token (using the TransformerLens framework, zero ablation of the blocks.0.hook_resid_post data set, the above loss function, and a “cut off” loss threshold of 0.08). Tab. 4 shows sample results.
 Token Average loss Conclusion
 0 .. 10 0.070 Impact is low. Calcs for these 11 tokens are unimportant
11 0.322 Loss is 4 x threshold. Calculations are important
12 0.497 Loss is 6 x threshold. Calculations are important
13 0.604 Loss is 7 x threshold. Calculations are important
14 0.921 Loss is 11 x threshold. Calculations are important
15 1.181 Loss is 14 x threshold. Calculations are important
16 1.021 Loss is 12 x threshold. Calculations are important
17 0.070 Impact is low. Calculations for this token is unimportant
Table 4: For 5 digit addition, for the 18 tokens / calculation rows, how abalting each token/row impacts the calculation loss.
This shows that all important calculations are completed while the model processes just 6 of the 18 tokens.
For deeper investigation, we created 1̃00 hand-curated test questions. They cover all question types (BA, MC1 and MS9) and all the answer digits (A5 .. A0) that these question types can occur in. These test cases were not used in model training. Example questions include:
  • make_a_question( _, _, 888, 11111, BASE_ADD_CASE)
  • make_a_question( _, _, 35000, 35000, USE_CARRY_1_CASE)
  • make_a_question( _, _, 15020, 45091, USE_CARRY_1_CASE)
  • make_a_question( _, _, 25, 79, SIMPLE_US9_CASE)
  • make_a_question( _, _, 41127, 10880, SIMPLE_US9_CASE)
  • make_a_question( _, _, 123, 877, CASCADE_US9_CASE)
  • make_a_question( _, _, 81818, 18182, CASCADE_US9_CASE)
The above experiment was repeated (with “cut off” loss threshold of 0.1) using these 1̃00 questions (and another 64 random questions). We analyzed the incorrect answers, grouping these failures by which answer digits were wrong. Tab. 5 shows sample results.
 Row Number of incorrect answers, grouped by incorrect digits
 11 ’Nyyyyy’: 47, ’NyNyyy’: 5, ’yNyyyy’: 7, ’yyNyyy’: 7, ’yyyNyy’: 1, …
12 ’yNyyyy’: 97, ’yNNyyy’: 7, ’NNyyyy’: 3, ’Nyyyyy’: 2, ’yyyNyy’: 1, …
13 ’yyNyyy’: 85, ’NyNyyy’: 3, ’yNNyyy’: 2, ’Nyyyyy’: 3, ’yNyyyy’: 5, …
14 ’yyyNyy’: 72, ’yyNNyy’: 4, ’NyyNyy’: 3, ’yNyNyy’: 3, ’Nyyyyy’: 2, …
15 ’yyyyNy’: 74, ’yyNyNy’: 3, ’Nyyyyy’: 2, ’NyyyNy’: 3, ’yNyyNy’: 3, …
16 ’yyyyyN’: 82, ’yyNyyN’: 2, ’NyyyyN’: 4, ’yNyyyN’: 4, ’yyNyyy’: 9, …
Table 5: For 5 digit addition, how ablating all the attention heads for a row impacts answer digit correctness. An ‘N’ means that answer digit was incorrect in the predicted answer. In each step, the failure count for the top grouping is > 6 times the next most common grouping - evidence that each step calculates one answer digit. There is a very steady progression through the rows with each successive row focusing on the next answer digit - this progression reflects numerically the “double staircase” attention pattern.
We also ablated one head at a time (using the TransformerLens framework, mean ablation of the blocks.0.attn.hook_z data set, the standard loss function, and a “cut off” loss threshold of 0.08) to understand which head(s) were key in calculating BA questions. Tab. 6 shows sample results.
The last row of Tab. 6 shows that digit A0 was calculated in step 16. This is one step before the model reveals A0 in step 17. The other rows in the table shows that this pattern holds true for all the other digits: each digit is calculated one step before the model reveals it.
 Ablated Head Average loss Conclusion
 0 0.016 Impact is low. Head is unimportant for BA
1 4.712 Loss is 50 x threshold. Head is important for BA
2 0.062 Impact is low. Head is unimportant for BA
Table 6: For 5 digit addition, when ablating heads, Head 1 is clearly key for the calculation of Base Add questions.